OpenStudy (anonymous):

-5(3-a)=-3(2-a) a=??

OpenStudy (alfie):

\[-15 + 5a = -6 + 3a \rightarrow 5a-3a = 15-6 \rightarrow 2a=9 \rightarrow a= 9/2\]